TIFF2PS(1) TIFF2PS(1) NAME tiff2ps−convertaTIFFimagetoPOSTSCRIPTTM SYNOPSIS tiff2ps[options]input.tif... DESCRIPTION tiff2psreadsTIFFimagesandwritesPOSTSCRIPTorEncapsulatedPOSTSCRIPT(EPS)onthe standardoutput.Bydefault,tiff2pswritesEncapsulatedPOSTSCRIPTforthefirstimageinthe specifiedTIFFimagefile. Bydefault,tiff2pswillgeneratePOSTSCRIPTthatfillsaprintedareaspecifiedbytheTIFFtagsin theinputfile.IfthefiledoesnotcontainXResolutionorYResolutiontags,thentheprintedarea issetaccordingtotheimagedimensions.The−wand−hoptions(seebelow)canbeusedtoset thedimensionsoftheprintedareaininches;overridinganyrelevantTIFFtags. ThePOSTSCRIPTgeneratedforRGB,palette,andCMYKimagesusesthecolorimageoperator. ThePOSTSCRIPTgeneratedforgreyscaleandbilevelimagesusestheimageoperator.Whenthe colorimageoperatorisused,POSTSCRIPTcodetoemulatethisoperatoronolderPOSTSCRIPT printersisalsogenerated.Notethatthisemulationcodecanbeveryslow. Colorimageswithassociatedalphadataarecompositedoverawhitebackground. OPTIONS −1GeneratePOSTSCRIPTLevelI(thedefault). −2GeneratePOSTSCRIPTLevelII. −aGenerateoutputforallIFDs(pages)intheinputfile. −dSettheinitialTIFFdirectorytothespecifieddirectorynumber.(NB:directoriesarenumberedstartingatzero.)Thisoptionisusefulforselectingindividualpagesinamulti-page (e.g.facsimile)file. −eForcethegenerationofEncapsulatedPOSTSCRIPT. −hSpecifytheverticalsizeoftheprintedarea(ininches). −oSettheinitialTIFFdirectorytotheIFDatthespecifiedfileoffset.Thisoptionisuseful forselectingthumbnailimagesandthelikewhicharehiddenusingtheSubIFDtag. −pForcethegenerationof(non-Encapsulated)POSTSCRIPT. −sGenerateoutputforasingleIFD(page)intheinputfile. −wSpecifythehorizontalsizeoftheprintedarea(ininches). −zWhengeneratingPOSTSCRIPTLevelII,dataisscaledsothatitdoesnotimageintothe deadzoneonapage(theoutermarginthattheprintingdeviceisunabletomark).This optionsuppressesthisbehaviour.WhenPOSTSCRIPTLevelIisgenerated,dataisimaged totheentireprintedpageandthisoptionhasnoaffect. EXAMPLES ThefollowinggeneratesPOSTSCRIPTLevelIIforallpagesofafacsimile: tiff2ps-a2fax.tif|lpr Notealsothatifyouhaveversion2.6.1ornewerofGhostscriptthenyoucanefficientlypreview facsimilegeneratedwiththeabovecommand. TogenerateEncapsulatedPOSTSCRIPTforatheimageatdirectory2ofanimageuse: tiff2ps-d1foo.tif (noticethatdirectoriesarenumberedstartingatzero.) BUGS BecausePOSTSCRIPTdoesnotsupportthenotionofacolormap,8-bitpaletteimagesproduce 24-bitPOSTSCRIPTimages.Thisconversionresultsinoutputthatissixtimesbiggerthanthe originalimageandwhichtakesalongtimetosendtoaprinteroveraserialline.Mattersare January27,1997 1 TIFF2PS(1) TIFF2PS(1) evenworsefor4-,2-,and1-bitpaletteimages. BUGS DoesnothandletiledimageswhengeneratingPSLevelIoutput. SEEALSO pal2rgb(1),tiffinfo(1),tiffcp(1),tiffgt(1),tiffmedian(1),tiff2bw(1),tiffsv(1),libtiff(3) January27,1997 2